# Proportionality for Constrained Public Decisions

### v1

@@ 5 Justified Representation with Agreement
Next, we show the representation guarantees can be achieved on instances with $k$-restrictive constraints via the use of modified version of $\text{MeCorA}_{\mathcal{C}}$. Moreover, we can drop the restriction to binary election instances that was key for the result of Theorem 14. In this $\text{MeCorA}_{\mathcal{C}}$ variant, we first partition the voter population into groups where members of each group agree on some set of issues. Then, for each group, its members may only pay to change some decisions as a collective and only on those issues that they agree on. Contrarily to $\text{MeCorA}_{\mathcal{C}}$, voter groups cannot pay to change some decisions if this leads to the group’s members gaining ‘too much’ satisfaction from the altered outcome (i.e., a voter group exceeding their proportional share of their agreed-upon issues, up to some additive factor $q$ that parameterises the rule).
@@
As with $\text{MeCorA}_{\mathcal{C}}$, voter groups shall pay to change the decisions of some issues during the rule’s execution. However, given the initial partition, during the run of Greedy $\text{MeCorA}_{\mathcal{C}}\text{-}q$, the voters in $N(T_{x})$ may only change decisions for the issues in $T_{x}$.
Moreover, if a voter group $N(T_{x})$ for some $x\in\{1,\ldots,p\}$ wishes to change some decisions at any moment during the process, this change does not lead to any voter in $N(T_{x})$ having satisfaction greater than $\nicefrac{{|N(T_{x})|}}{{n}}\cdot|T_{x}|-q$ with the updated outcome. Besides these two differences, the rule works exactly as $\text{MeCorA}_{\mathcal{C}}$.
Now, we can show the following for Greedy $\text{MeCorA}_{\mathcal{C}}\text{-}q$ working on a $k$-restrictive constraint. For this result, we require the additional assumption that voter ballots represent feasible outcomes in $\mathcal{C}$.
@@
In this case, the voters in $N^{\prime}$ would have at least $\nicefrac{{nmk}}{{|T|}}$ in collective funds, so it follows that each distinct $(S,\boldsymbol{w})$-deviation available to $N^{\prime}$ must cost at least $\nicefrac{{nmk}}{{|T|}}$. As $N^{\prime}$ is $T$-agreeing, it must be that $N^{\prime}$ has at least a $\nicefrac{{(|T|-\ell+1)}}{{k}}$ many $(S,\boldsymbol{w})$-deviations due to $\mathcal{C}$ being $k$-restrictive and as the voters’ ballots are consistent with $\mathcal{C}$.
Now, consider the case where some voter in $N^{\prime}$ contributed more than $\nicefrac{{m}}{{(\ell+k-1)}}$ to change some decisions. The first time that this occurred, the change of decisions did not lead to any voter in $N^{\prime}$ obtaining a satisfaction greater than $\ell=\nicefrac{{|N^{\prime}|}}{{n}}\cdot|T|-k+1$ (otherwise the rule would not allow these voters to pay for the changes). Thus, each voter in $N^{\prime}$ must have contributed to at most $\ell-1$ issues before this moment. From the reasoning above, it must hold that in this moment, each voter held at least $\nicefrac{{nmk}}{{|N^{\prime}||T|}}$ in funds with there being at least $\nicefrac{{(|T|-\ell+1)}}{{k}}$ feasible deviations available to $N^{\prime}$ and each such deviation costing at least $\nicefrac{{nmk}}{{|T|}}$. So in both cases, for the $(S,\boldsymbol{w})$-deviations that are present in $T$ that voters in $N^{\prime}$ wish to make, outcome $\boldsymbol{w}$’s decisions must have been paid for by voters within the remaining voter population $N\setminus N^{\prime}$. And so, these decisions must have cost the voters in $N\setminus N^{\prime}$ at least:
@@
$>\frac{nm}{|T|}\cdot\bigg{(}\frac{n|T|-|N^{\prime}||T|}{n}\bigg{)}=m(n-|N^{\prime}|).$
@@
We now offer another way towards producing proportional outcomes when using $k$-restrictive constraints. It is a constrained adaptation of the Local Search Proportional Approval Voting (LS PAV) rule from the MWV literature that is a polynomial-time computable rule that is known to satisfy EJR (Aziz et al., 2017). In the MWV setting, the rule begins with an arbitrary committee of some fixed size $k$ and in iterations, searches for any swaps between committee members and non-selected candidates that brings about an increase of the PAV score by at least $\nicefrac{{n}}{{k^{2}}}$. In our model, the PAV score of some feasible outcome $\boldsymbol{w}\in\mathcal{C}$ is defined to be $\text{PAV}(\boldsymbol{w})=\sum_{i\in N}\sum_{t=1}^{u_{i}(\boldsymbol{w})}\nicefrac{{1}}{{t}}$. We can then lift Local Search PAV to our setting with constraints.
Definition 13 (Local Search $\text{PAV}_{\mathcal{C}}$, LS $\text{PAV}_{\mathcal{C}}$).
Beginning with an arbitrary outcome $\boldsymbol{w}\in\mathcal{C}$ as the current winning outcome, the rule looks for all possible deviations. If there exists an ($S$,$\boldsymbol{w}$)-deviation for some voter group to some outcome $\boldsymbol{w}^{\prime}\in\mathcal{C}$ such that $PAV(\boldsymbol{w}^{\prime})-PAV(\boldsymbol{w})\geqslant\nicefrac{{n}}{{m^{2}}}$, i.e., the new outcome $\boldsymbol{w}^{\prime}$ yields a PAV score that is at least $\nicefrac{{n}}{{m^{2}}}$ higher than that of $\boldsymbol{w}$, then the rule sets $\boldsymbol{w}^{\prime}$ as the current winning outcome. The rule terminates once there exists no deviation that improves on the PAV score of the current winning outcome by at least $\nicefrac{{n}}{{m^{2}}}$.
As there is a maximum obtainable PAV score, LS $\text{PAV}_{\mathcal{C}}$ is guaranteed to terminate. The question is how long this rule takes to return an outcome when we have to take $k$-restrictive constraints into account.
@@
For elections instances where $\mathcal{C}$ is $k$-restrictive (where $k$ is a fixed constant), LS $\text{PAV}_{\mathcal{C}}$ terminates in polynomial time.
@@
Off the back of this positive computational result, we present the degree to which LS $\text{PAV}_{\mathcal{C}}$ provides proportional outcomes with regards to the $\alpha\text{-}{\text{agrEJR}_{\mathcal{C}}}\text{-}\beta$ axiom.
@@
Now, so there are $|T|-(\ell-1)=|T|-\ell+1$ issues that all voters in $N^{\prime}$ agree on but they disagree with outcome $\boldsymbol{w}$’s decisions on these issues. Since we assume the constraint is $k$-restrictive, then for each of these $|T|-\ell+1$ issues, they fix at most $k-1$ other issues and thus, there are at least $\nicefrac{{(|T|-\ell+1)}}{{k}}$ feasible deviations that can be made by LS-$\text{PAV}_{\mathcal{C}}$ amongst these issues. For the voters in $N^{\prime}$, we now consider the minimal increase in PAV score that may occur from these possible deviations by LS-$\text{PAV}_{\mathcal{C}}$. For each such deviation, we increase their satisfaction by at least $k$ and thus, for a voter $i\in N^{\prime}$, we increase the PAV score by $\sum_{t=1}^{k}\nicefrac{{1}}{{(r_{i}+t)}}$. Since for each voter $i\in N^{\prime}$ we have $r_{i}\leqslant\ell-1$, and as there are at least $\nicefrac{{(|T|-\ell+1)}}{{k}}$ feasible deviations in $T$, it follows that we add at least the following to the PAV score:
@@
$>\frac{n(k+1)}{2}-|N^{\prime}|+\frac{n(k+1)}{|T|}\geqslant\frac{k+1}{2}\cdot(n-|N^{\prime}|)+\frac{n(k+1)}{|T|}.$
@@ 6 Proportionality via Priceability
With this section, we offer an alternative to the justified-representation-like interpretation of proportional representation, and this is through the notion of priceability (Lackner and Skowron, 2023). Recent work has shown the promise of this market-based approach for a general social choice model (Masařík et al., 2023) and the sequential choice model (Chandak et al., 2024). We look to employ it for constrained public decisions (albeit looking at a weaker priceability axiom than the axiom that Masařík et al. (2023) studied).
@@
[…]
@@
[…]
@@
Consider a priceable outcome $\boldsymbol{w}$ with price system $\textbf{ps}=(\{p_{i}\}_{i\in N},\{\pi_{(a_{t},d)}\}_{(a_{t},d)\in H})$ where $H=\bigcup_{a_{t}\in\mathcal{I}}\{(a_{t},d)\mid d\in D_{t}\}$. Then, for every $T$-cohesive group of voters $N^{\prime}\subseteq N$ for some $T\subseteq\mathcal{I}$ with an $(S,\boldsymbol{w})$-deviation for some $S\subseteq T$, it holds that:
@@
Take a priceable outcome $\boldsymbol{w}$ and consider a $T$-cohesive group of voters $N^{\prime}$. Suppose that $\sum_{i\in N^{\prime}}u_{i}(\boldsymbol{w})<\nicefrac{{n}}{{q}}\cdot|T|-|S|$ where $q=\max\{\pi_{(a_{t},w_{t})}\}_{a_{t}\in S}$. As a group, the voters $N^{\prime}$ have a budget of $m|N^{\prime}|$. Now, the voters in $N^{\prime}$ collectively contributed to at most $\nicefrac{{n}}{{q}}\cdot|T|-|S|-1$ decisions in outcome $\boldsymbol{w}$, and for each decision, the price was at most $q$ (as $q$ is the the price system’s maximal price). So, we have that voter group $N^{\prime}$ has at least the following in leftover budget:
@@
[…]
@@
[…]
This is a positive result that, combined with that of Proposition 18, gives us a rule that always returns us priceable outcomes for any election instance.

### v2

@@ 5 Justified Representation with Agreement
Example 11.
Suppose we have a election instance with three issues $\mathcal{I}=\{a_{1},a_{2},a_{3}\}$ with domains $D_{t}=\{d_{1},d_{2},d_{3}\}$ for all $t\in\{1,2,3,4\}$. Then take the constraint to be $\mathcal{C}=\{(d_{1},d_{1},d_{1}),(d_{1},d_{1},d_{2}),(d_{1},d_{1},d_{3}),(d_{2},d_{2},d_{1})\}$. Then, suppose there are four voters $N=\{1,2,3\}$ with ballots $\boldsymbol{b}_{1}=(d_{1},d_{1},d_{1})$, $\boldsymbol{b}_{2}=(d_{1},d_{1},d_{2})$, $\boldsymbol{b}_{3}=(d_{2},d_{2},d_{2})$ and $\boldsymbol{b}_{4}=(d_{3},d_{3},d_{3})$.
Assume that the rule begins with the current outcome $\boldsymbol{w}=(d_{2},d_{2},d_{1})$ where the price for each issue is 0 and the individual voter budgets are $m=3$. As voters $\{1,2\}$ have an $(S,\boldsymbol{w},\mathcal{C})$-deviation for the first two issues $S=\{a_{1},a_{2}\}$ and they each have their entire budgets available, they can pay to update the current outcome to $\boldsymbol{w}^{*}=(d_{1},d_{1},d_{1})$. This update raises the price of each of issues $\{a_{1},a_{2}\}$ by $\epsilon$. In the next instance, there are feasible deviations on a single issue that are available for the groups $\{2,3\}$ (to outcome $(d_{1},d_{1},d_{2})$), $\{3\}$ (to outcome $(d_{2},d_{2},d_{1})$) and $\{4\}$ (to outcome $(d_{1},d_{1},d_{3})$). All of these deviations are affordable however, the rule favours the voter group $\{2,3\}$ as it is cheaper per voter. Thus, the current outcome is updated to $\boldsymbol{w}^{**}=(d_{1},d_{1},d_{2})$ and the price to update issue $a_{3}$ is raised to $\epsilon$. Finally, suppose that at some future instance, the current outcome is again $\boldsymbol{w}=(d_{1},d_{1},d_{2})$ and suppose that the price of issue is at least $3$ (this is possible as each voter has a budget of $3$). Then, although voter $4$ has an $(S,\boldsymbol{w},\mathcal{C})$-deviation to outcome $(d_{1},d_{1},d_{3})$ and all of their budget (as they have zero satisfaction), they cannot afford to raise the price of issue $a_{3}$ so cannot update the outcome.
Our definition of $\text{MeCorA}_{\mathcal{C}}$ will be useful in Section 6, while now we need to refine its definition further to study its justified representation guarantees. We first partition the voter population into groups where members of each group agree on some set of issues. Then, for each group, its members may only pay to change some decisions as a collective and only on those issues that they agree on. Contrarily to $\text{MeCorA}_{\mathcal{C}}$, voter groups cannot pay to change some decisions if this leads to the group’s members gaining “too much” satisfaction from the altered outcome (i.e., a voter group exceeding their proportional share of their agreed-upon issues, up to some additive factor $q$ that parameterises the rule).
@@
As with $\text{MeCorA}_{\mathcal{C}}$, voter groups pay to change the decisions of some issues during the rule’s execution. However, given the initial partition, during the run of Greedy $\text{MeCorA}_{\mathcal{C}}\text{-}q$, the voters in $N(T_{x})$ may only change decisions for the issues in $T_{x}$. Moreover, if a voter group $N(T_{x})$ for some $x\in\{1,\ldots,p\}$ wishes to change some decisions at any moment during the process, this change does not lead to any voter in $N(T_{x})$ having satisfaction greater than $\nicefrac{{|N(T_{x})|}}{{n}}\cdot|T_{x}|-q$ with the updated outcome for some non-negative integer $q$. Besides these two differences, the rule works exactly as $\text{MeCorA}_{\mathcal{C}}$.
Consider the following example that demonstrates how Greedy $\text{MeCorA}_{\mathcal{C}}\text{-}q$ works.
Example 12.
Take an election instance with five issues $\mathcal{I}=\{a_{1},a_{2},a_{3},a_{4},a_{5}\}$ with domains $D_{t}=\{d_{1},d_{2},d_{3}\}$ for all $t\in\{1,2,3,4,5\}$. Then, suppose there are five voters $N=\{1,2,3,4,5\}$ with ballots $\boldsymbol{b}_{1}=(d_{1},d_{1},d_{1},d_{1},d_{1})$, $\boldsymbol{b}_{2}=(d_{1},d_{1},d_{1},d_{1},d_{1})$, $\boldsymbol{b}_{3}=(d_{2},d_{2},d_{2},d_{1},d_{1})$, $\boldsymbol{b}_{4}=(d_{2},d_{2},d_{2},d_{2},d_{2})$ and $\boldsymbol{b}_{5}=(d_{3},d_{3},d_{3},d_{3},d_{3})$. The rule first partitions the voters into three groups $\{1,2\}$, $\{3,4\}$ and $\{4\}$ where the voter groups can only update the decisions of issues in $\{a_{1},a_{2},a_{3},a_{4},a_{5}\}$, $\{a_{1},a_{2},a_{3},a_{4}\}$ and $\{a_{1},a_{2},a_{3},a_{4},a_{5}\}$, respectively. The only other difference to $\text{MeCorA}_{\mathcal{C}}$ is that for the voter groups $\{1,2\}$, $\{3,4\}$ and $\{4\}$, the satisfaction garnered from the new outcome in any deviating update cannot exceed a certain amount. Suppose that $q=0$ for Greedy $\text{MeCorA}_{\mathcal{C}}\text{-}q$. For voter groups $\{1,2\}$, $\{3,4\}$ and $\{4\}$, this satisfaction limit is set to $\nicefrac{{2}}{{5}}\cdot 5=2$, $\nicefrac{{2}}{{5}}\cdot 3=\nicefrac{{6}}{{5}}$ and $\nicefrac{{1}}{{5}}\cdot 5=1$, respectively.
We can now show our main result for Greedy $\text{MeCorA}_{\mathcal{C}}\text{-}q$ on a $k$-restrictive constraint. For this result, we require that voter ballots are feasible outcomes in $\mathcal{C}$, but we do not need to assume that we are in a binary election instance.
@@
In this case, the voters in $N^{\prime}$ would have at least $\nicefrac{{nmk}}{{|T|}}$ in collective funds, so it follows that each distinct $(S,\boldsymbol{w},\mathcal{C})$-deviation available to $N^{\prime}$ must cost at least $\nicefrac{{nmk}}{{|T|}}$. As $N^{\prime}$ is $T$-agreeing, it must be that $N^{\prime}$ has at least a $\nicefrac{{(|T|-\ell+1)}}{{k}}$ many $(S,\boldsymbol{w},\mathcal{C})$-deviations due to $\mathcal{C}$ being $k$-restrictive and as the voters’ ballots are consistent with $\mathcal{C}$.
Now, consider the case where some voter in $N^{\prime}$ contributed more than $\nicefrac{{m}}{{(\ell+k-1)}}$ to change some decisions. The first time that this occurred, the change of decisions did not lead to any voter in $N^{\prime}$ obtaining a satisfaction greater than $\ell=\nicefrac{{|N^{\prime}|}}{{n}}\cdot|T|-k+1$ (otherwise the rule would not allow these voters to pay for the changes). Thus, each voter in $N^{\prime}$ must have contributed to at most $\ell-1$ issues before this moment. From the reasoning above, it must hold that in this moment, each voter held at least $\nicefrac{{nmk}}{{|N^{\prime}||T|}}$ in funds with there being at least $\nicefrac{{(|T|-\ell+1)}}{{k}}$ feasible deviations available to $N^{\prime}$ and each such deviation costing at least $\nicefrac{{nmk}}{{|T|}}$. So in both cases, for the $(S,\boldsymbol{w},\mathcal{C})$-deviations that are present in $T$ that voters in $N^{\prime}$ wish to make, outcome $\boldsymbol{w}$’s decisions must have been paid for by voters within the remaining voter population $N\setminus N^{\prime}$. And so, these decisions must have cost the voters in $N\setminus N^{\prime}$ at least:
@@
$\frac{nm}{|T|}\cdot\bigg(\frac{n|T|-|N^{\prime}||T|}{n}\bigg)=m(n-|N^{\prime}|).$
@@
We now explore another direction towards producing proportional outcomes on $k$-restrictive constraints. We define a constraint-aware version of the Local Search Proportional Approval Voting (LS-PAV) rule, which is a polynomial-time computable rule that is known to satisfy EJR (Aziz et al., 2017). In the committee elections setting, the rule begins with an arbitrary committee of some fixed size $k$, and improves it in iterations by searching for any swaps between committee members and non-selected candidates that brings about an increase of the PAV score by at least $\nicefrac{{n}}{{k^{2}}}$. In our model, the PAV score of some feasible outcome $\boldsymbol{w}\in\mathcal{C}$ is defined to be $\text{PAV}(\boldsymbol{w})=\sum_{i\in N}\sum_{t=1}^{u_{i}(\boldsymbol{w})}\nicefrac{{1}}{{t}}$.
Definition 14 (Local Search $\text{PAV}_{\mathcal{C}}$, LS-$\text{PAV}_{\mathcal{C}}$).
[…]
Example 13.
[…]
As there is a maximum obtainable PAV score, LS-$\text{PAV}_{\mathcal{C}}$ is guaranteed to terminate. The question is how long this rule takes to return an outcome when we have to take $k$-restrictive constraints into account.
@@
For elections instances where $\mathcal{C}$ is $k$-restrictive (where $k$ is a fixed constant), LS-$\text{PAV}_{\mathcal{C}}$ terminates in polynomial time.
@@
Off the back of this positive computational result, we present the degree to which LS-$\text{PAV}_{\mathcal{C}}$ provides proportional outcomes with regards to the $\alpha\text{-}{\text{agrEJR}_{\mathcal{C}}}\text{-}\beta$ axiom.
@@
Now, so there are $|T|-(\ell-1)=|T|-\ell+1$ issues that all voters in $N^{\prime}$ agree on but they disagree with outcome $\boldsymbol{w}$’s decisions on these issues. Since we assume the constraint is $k$-restrictive, then for each of these $|T|-\ell+1$ issues, they fix at most k1k-1 other issues and thus, there are at least $\nicefrac{{(|T|-\ell+1)}}{{k}}$ feasible deviations that can be made by LS-$\text{PAV}_{\mathcal{C}}$ amongst these issues. For the voters in $N^{\prime}$, we now consider the minimal increase in PAV score that may occur from these possible deviations by LS-$\text{PAV}_{\mathcal{C}}$. For each such deviation, we increase their satisfaction by at least $k$ and thus, for a voter $i\in N^{\prime}$, we increase the PAV score by $\sum_{t=1}^{k}\nicefrac{{1}}{{(r_{i}+t)}}$. Since for each voter $i\in N^{\prime}$ we have $r_{i}\leqslant\ell-1$, and as there are at least $\nicefrac{{(|T|-\ell+1)}}{{k}}$ feasible deviations in $T$, it follows that we add at least the following to the PAV score:
@@
$\frac{n(k+1)}{2}-|N^{\prime}|+\frac{n(k+1)}{|T|}\geqslant\frac{k+1}{2}\cdot(n-|N^{\prime}|)+\frac{n(k+1)}{|T|}.$
@@ 6 Proportionality via Priceability
In this section we propose an alternative to the interpretation of proportional representation as justified representation through the notion of priceability (Peters and Skowron, 2020; Lackner and Skowron, 2023). Recent work has shown the promise of this market-based approach for a general social choice model (Masarík et al., 2024) and the sequential choice model (Chandak et al., 2024). We propose the following version for constrained public decisions albeit looking at a weaker priceability axiom than that of Masarík et al. (2024) as they use an adaptation of stable priceability (Peters et al., 2021a) that is not always satisfiable (Masarík et al., 2024) (see Section $6.2$ in their paper).
@@
  1. $(\rm P1):$ For all $a_{t}\in\mathcal{I}$ and $d\in D_{t}$, it holds that if $d\neq\boldsymbol{b}_{i,t}$ we have $p_{i}(a_{t},d)=0$, for every $i\in N$.
  2. $(\rm P2):$ $\sum_{(a_{t},d)\in H}p_{i}(a_{t},d)\leqslant m$ for every $i\in N$ where it holds that $H=\bigcup_{a_{t}\in\mathcal{I}}\{(a_{t},d)\mid d\in D_{t}\}$.
  3. $(\rm P3):$ $\sum_{i\in V}p_{i}(a_{t},d)=\pi_{(a_{t},w_{t})}$ for every $a_{t}\in\mathcal{I}$.
  4. $(\rm P4):$ $\sum_{i\in V}p_{i}(a_{t},d)=0$ for every $a_{t}\in\mathcal{I}$ and every $d\neq w_{t}\in D_{t}$.
  5. $(\rm P5):$ There exists no group of voters $N^{\prime}$ with an $(S,\boldsymbol{w},\mathcal{C})$-deviation for some $S\subseteq\mathcal{I}$, such that for each $a_{t}\in S$:
@@
Take four issues $\mathcal{I}=\{a_{1},a_{2},a_{3},a_{4}\}$ in a binary election instance and a constraint $\mathcal{C}=\{(1,1,1,1),(1,1,0,0)\}$. Suppose there are two voters with ballots $\boldsymbol{b}_{1}=(1,1,1,1)$ and $\boldsymbol{b}_{2}=(0,0,0,0)$. Note that outcome $\boldsymbol{w}=(1,1,1,1)$ is not priceable as any price system where voter $1$ does not exceed her budget would have voter $2$ having enough in leftover budget to cause a violation of condition $(\rm P5)$ (with her entire budget being leftover, she can afford more than the price of the $(S,\boldsymbol{w},\mathcal{C})$-deviation to outcome $\boldsymbol{w}$). On the other hand, $\boldsymbol{w}^{\prime}=(1,1,0,0)$ is priceable where we set the price of this outcome’s decisions to $1$.
@@
Consider a priceable outcome $\boldsymbol{w}$ with price system $\textbf{ps}=(\{p_{i}\}_{i\in N},\{\pi_{(a_{t},d)}\}_{(a_{t},d)\in H})$ where $H=\bigcup_{a_{t}\in\mathcal{I}}\{(a_{t},d)\mid d\in D_{t}\}$. Then, for every $T$-cohesive group of voters $N^{\prime}\subseteq N$ for some $T\subseteq\mathcal{I}$ with an $(S,\boldsymbol{w},\mathcal{C})$-deviation for some non-empty $S\subseteq T$, it holds that:
@@
Take a priceable outcome $\boldsymbol{w}$ and consider a $T$-cohesive group of voters $N^{\prime}$. Suppose that $\sum_{i\in N^{\prime}}u_{i}(\boldsymbol{w})<\nicefrac{{n}}{{q}}\cdot|T|-|S|$ where $q=\max\{\pi_{(a_{t},w_{t})}\}_{a_{t}\in S}$. As a group, the voters $N^{\prime}$ have a budget of $m|N^{\prime}|$. Now, the voters in $N^{\prime}$ collectively contributed to at most $\nicefrac{{n}}{{q}}\cdot|T|-|S|-1$ decisions in outcome $\boldsymbol{w}$, and for each decision, the price was at most $q$ (as $q$ is the price system’s maximal price). So, we have that voter group $N^{\prime}$ has at least the following in leftover budget:
@@
Note we made use of the fact that $N^{\prime}$ is $T$-cohesive. Thus, we know that $N^{\prime}$ has strictly more than $q|S|$ in funds and for each issue in $a_{t}\in S$, holds more than in funds than $q\geqslant\pi_{(a_{t},w_{t})}$. This presents a violation of condition $\rm P5$} of priceability. Hence, voter group $N^{\prime}$ cannot exist. ∎
@@
Let $\boldsymbol{w}=(w_{1},\ldots,w_{m})$ be the outcome returned by $\text{MeCorA}_{\mathcal{C}}$. We define the following price system ps: For each issue $a_{t}\in\mathcal{I}$, fix the prices $\pi_{(a_{t},w_{t})}=\pi_{(a_{t},d)}=\lambda_{t}$ for all $d\neq w_{t}\in D_{t}$ where $\lambda_{t}$ is issue $a_{t}$’s last $\text{MeCorA}_{\mathcal{C}}$ price (before being set to $\infty$) prior to the rule’s termination. Fix the payment functions $p_{i}$ for each voter to the money they spent to end the execution of $\text{MeCorA}_{\mathcal{C}}$. Observe that the priceability conditions $(\rm P1)$-$(\rm P4)$ clearly hold: since we have that, to end $\text{MeCorA}_{\mathcal{C}}$’s run, voters do not pay for decisions that $(i)$ they do not agree with (condition $(\rm P1)$) and $(ii)$ are not made by outcome $\boldsymbol{w}$ (condition $(\rm P4)$); $\text{MeCorA}_{\mathcal{C}}$ limits each voter a budget of $m$ (condition $(\rm P2)$) $(\rm P2)$; and the sum of payments for decisions made by outcome $\boldsymbol{w}$ will equal exactly $\pi_{(a_{t},w_{t})}=\lambda_{t}$ (condition $(\rm P3)$). Now, for condition $(\rm P5)$, note that if such a group of voters $N^{\prime}$ existed for some set of issues $S$, then $\text{MeCorA}_{\mathcal{C}}$ would not have terminated as this group of voters could have changed the decisions of these issues in $S$ while increasing each issues’ prices. ∎
This is a positive result that, combined with that of Proposition 18, gives a rule that always returns us priceable outcomes for any election instance.